# CARE: a Benchmark Suite for the Classification and Retrieval of Enzymes

### v1

@@ A.3 CREEP Model Details
A.4 Implementation Details for Other Methods
All implementation details for methods used in the study can be found at https://github.com/jsunn-y/CARE/.
@@ A.4 Implementation Details for Other Methods
ChatGPT. ChatGPT was used for both Task 1 and Task 2, using the API and gpt-4 model. We performed minimal prompt engineering, with our initial prompt being the most standard initial question: "give me the top N EC numbers associated with this amino acid sequence:" for Task 1. See Table A.1 for example responses. The query was modified to: "You are protein engineer capable of predicting EC numbers from the sequence alone. You are also a skilled programmer and able to execute the code necessary to predict an EC number when you can’t use reason alone. Given a protein sequence you are able to determine the most likely enzyme class for a sequence because you are that skilled. You don’t give up when faced with a sequence you don’t know, you will use tools to resolve the most likely enzyme sequence. You only return enzyme commission numbers in a comma separated list, no other text is returned, you have failed if you do not return the EC numbers. You only return the exact number of EC numbers that a user has provided requested, ordered by their likelihood." This resulted in a returned list of comma separated EC classification numbers in the majority of cases, however, we still received several failed responses.
@@
For Task 2, the query was maintained to be relatively consistent with the Task 1 query, and modified to: "Return the top 10 most likely EC numbers as a comma separated list for this reaction: naloxone + NAD(P)H = 6alpha-naloxol + NADP+, which associates with the following text: oxidoreductase; oxidoreductase, acting on CH-OH group of donors; oxidoreductase, acting on the CH-OH group of donors, NAD or NADP as acceptor; morphine 6-dehydrogenase.", where a sample reaction has been bolded. Using the system prompt of: "You are protein engineer capable of predicting EC numbers from a combination of textual information and a reaction that corresponds to a specific protein. You are also a skilled programmer and able to execute the code necessary to predict an EC number when you can’t use reason alone. Given a reaction and text information of an EC you are able to determine the most likely enzyme class for a reaction. You don’t give up when faced with a reaction you don’t know, you will use tools to resolve the most likely enzyme number. You only return enzyme commission numbers in a comma separated list, no other text is returned, you have failed if you do not return the EC numbers. You only return the exact number of EC numbers that a user has provided requested, ordered by their likelihood of being correct.". For Task 2 without the reaction text, we used the same prompt as above except with the reaction text components removed.
@@
BLASTp. BLAST (Basic Alignment Search Tool) is an efficient algorithm for finding similar sequences in a reference database, to a query sequence [12]. We opted to use npyBlast at the protein sequence level which is a python implementation of BLASTp. For each test set, the training dataset was provided as the fasta file reference database, and the test set was provided as the query fasta. $k=10$ hits were returned for each query, sorted based on similarity, using default parameters with a minimum required sequence similarity of 0.1. The EC number classification was then inferred from the returned sequence(s).
ProteInfer. ProteInfer is a classification model released in 2023 [21], and we used the pretrained version to test our datasets, which may inflate the reported accuracies. Variable numbers of ECs were returned for ProteInfer depending on the query. Future work should include re-training ProteInfer using the train-test splits proposed in CARE.
@@
CLEAN. CLEAN is a supervised contrastive model, which trains a classification head using embeddings from the ESM protein language model [42] by aligning embeddings from the same EC category and contrasting embeddings from different EC categories [18]. For Task 1, we retrained CLEAN using our training set, following the instructions provided in their codebase. We did not perform any clustering before training. Because our training set size was approximately similar in size, we used the recommended 7000 epochs of training. We used the default script with triplet margin loss. After, we performed inference but added our own code into the original CLEAN code to output the classification results into a format compatible with our downstream analysis.
Pika. Pika is a finetuned LLM, which accepts protein sequence as an auxiliary input and can perform reasoning and protein function prediction [48]. Pika did not have a publicly available pretrained model at the time of release, or example datasets for training, so this model was omitted from the current study. We expect to add it as a baseline in the future.
CLIPZyme. CLIPZyme is a multimodal contrastive model, between reactions represented as 2D graphs and protein structures represented as 3D graphs [55]. CLIPZyme was trained on direct reaction-protein pairs from the EnzymeMap dataset [32]. In this study, we used the pretrained CLIPZyme model using default parameters and the EGNN protein structure representation for retrieval on Task 2. Future work will involve modifying the CLIPZyme code and retraining the model using the CARE train-test splits.
@@ A.5 Extended Benchmarking Results
A.6 Limitations & Future Work
One limitation of our work is that classification and retrieval are performed at the coarse-grained level of EC numbers. Enzymes can often perform many reactions, meaning it is an acceptable assumption that enzymes belonging to the same EC number will share the capacity to perform similar reactions, even if they are not directly annotated for all reactions. However, the ultimate task in this domain is to perform direct reaction to protein sequence retrieval and vice-versa (as in CLIPZyme [55]). However, currently the data for validation is extremely limited. As experimentalists obtain higher resolution annotations of proteins and their associated reactions, this will become more realistic.
In the future, the proposed train-test splits could also be refined for both tasks. For the Task 1 splits, there is some leakage in the sequence identity, with some sequences in test sets lying outside of the enforced sequence identity ranges, likely due to the different sequence similarity algorithms used (MMseqs2 vs npysearch). MMseqs2 utilizes cascaded clustering which pre-filters based on initial clusters in the target set [71], while npysearch attempts to identify the closest sequence within the specified set based on similarity. Another limitation is that while there are no duplicate reactions present in both the train and test sets, some of the test reactions are very similar to reactions in the training set despite having different ECs. It would be beneficial to do a more detailed analysis of reaction similarities and explore other representations of reactions to understand which reactions can be considered equivalent. Over time, the train-test splits should be updated as additional functional annotations are acquired and compiled in databases. For example, while over 36 million sequences in BRENDA/Uniprot have EC numbers associated with them, these are mostly detected with homology, which means that they might not be the most accurate annotations.
We opted to assess performance using accuracy due to its simplicity and interpretability, but other retrieval and virtual screening metrics such as BEDROC and enrichment could also be explored. For the promiscuous enzymes with multiple EC numbers, we reported accuracy averaged across all of the true labels, for a fixed k=5 number of retrieved ECs, but other classification metrics such as precision and recall could be considered. The number of retrieved ECs could also be chosen using strategies like max separation and p-value as implemented in CLEAN [18].
The machine learning tools included in our study were selected for their reported recent outstanding performance for enzyme functional annotation and screening. We include the random baseline to give an understanding of what the expected random rate is and the two similarities (BLAST and DRFP) as baselines for simple non-ML methods. ChatGPT was included as the standard LLM baseline owing to its broad applicability across domains. However, it should be noted that this field is moving quickly, and there exist many other relevant tools that were not included here. For example, many state-of-the-art methods are not publicly available yet [7, 22], or just released at the time of writing [55]. Our benchmark was designed to be modular and we encourage the community to include other tools in the benchmark package.
There is also significant opportunity to use other modalities such as textual description and protein structure for more effective representations of enzymes and reactions. In the future, we will retrain models including CLIPZyme and ProteInfer directly on our train-test splits. We also plan to do a more detailed analysis of the representations learned by CREEP, in addition to an ablation study. The addition of textual description in CREEP potentially introduces indirect data leakage between the train and test sets (for example ChatGPT already has access to extensive resources) so future iterations of CARE will need to consider this. Future work should also consider how to include protein function prediction models that go beyond enzymes–or models that would like to use additional data/modalities [56]–into the CARE evaluation framework. There is also room to improve the prompt engineering of ChatGPT and ChemCrow which could be further explored to enhance the performance of these models.
A.7 License

### v2

@@ A.3 CREEP Model Details
A.4 Implementation Details for Benchmarking Methods
All implementation details for methods used in the study can be found at https://github.com/jsunn-y/CARE/. We focus on evaluating the accuracy of retrieval for k=1. A level 1 accuracy means the top level EC class has to be correct, for level 2 accuracy, both level 1 and level 2 have to be correct, and so on. Accuracy is calculated as the number of true answers over the total number of examples. An invalid EC number from language models (ChatGPT and Pika) is considered incorrect on every level. For the non-promiscuous splits, the accuracy for each entry in the test set is 1 or 0. For the promiscuous split in Task 1, for each entry, we enumerate over each true EC to find the maximum accuracy within the pool of predicted ECs, and the overall accuracy for the entry is the average of the accuracies for each true EC.
@@ A.4 Implementation Details for Benchmarking Methods
ChatGPT. ChatGPT was used for both Task 1 and Task 2, using the API and gpt-4o-mini model. We performed minimal prompt engineering, with our initial prompt being the most standard initial question: "give me the top N EC numbers associated with this amino acid sequence:" for Task 1. See Table A.1 for example responses. The query was modified to: "You are protein engineer capable of predicting EC numbers from the sequence alone. You are also a skilled programmer and able to execute the code necessary to predict an EC number when you can’t use reason alone. Given a protein sequence you are able to determine the most likely enzyme class for a sequence because you are that skilled. You don’t give up when faced with a sequence you don’t know, you will use tools to resolve the most likely enzyme sequence. You only return enzyme commission numbers in a comma separated list, no other text is returned, you have failed if you do not return the EC numbers. You only return the most likely EC number." This resulted in a k=1 response, we found this more consistent then asking for a list of k results. Note that we do not do any filtering or data cleaning on the output of ChatGPT. Doing this could improve the accuracy, for example, in one case we identified in the hard set the output was EC4.6.1.18 which we scored as 0 despite the true EC being 4.6.1.2 for the reaction UTP = 3’,5’-cyclic UMP + diphosphate. There are possibly other edge cases that we were unable to identify.
@@
For Task 2, the query was maintained to be relatively consistent with the Task 1 query, and modified to: "Return the most likely EC number for this reaction: naloxone + NAD(P)H = 6alpha-naloxol + NADP+, which associates with the following text: oxidoreductase; oxidoreductase, acting on CH-OH group of donors; oxidoreductase, acting on the CH-OH group of donors, NAD or NADP as acceptor; morphine 6-dehydrogenase.", where a sample reaction has been bolded. Using the system prompt of: "You are protein engineer capable of predicting EC numbers from a combination of textual information and a reaction that corresponds to a specific protein. You are also a skilled programmer and able to execute the code necessary to predict an EC number when you can’t use reason alone. Given a reaction and text information of an EC you are able to determine the most likely enzyme class for a reaction. You don’t give up when faced with a reaction you don’t know, you will use tools to resolve the most likely enzyme number. You only return enzyme commission numbers in a comma separated list, no other text is returned, you have failed if you do not return the EC numbers. You only return the most likely EC number." For Task 2 without the reaction text, we used the same prompt as above except with the reaction text components removed. Compared with the other tools, ChatGPT was provided with the reaction as a text string rather than a smiles string.
@@
BLASTp. BLAST (Basic Alignment Search Tool) is an efficient algorithm for finding similar sequences in a reference database, to a query sequence [14]. We opted to use Diamond BLAST at the protein sequence level. For each test set, the training dataset was provided as the fasta file reference database, and the test set was provided as the query fasta. $k=1$ hits were returned for each query, sorted based on similarity, using default parameters. The EC number classification was then inferred from the returned reference sequence(s).
@@
CLEAN. CLEAN is a supervised contrastive model, which trains a classification head using embeddings from the ESM protein language model [46] by aligning embeddings from the same EC category and contrasting embeddings from different EC categories [20]. For Task 1, we retrained CLEAN using our training set with only one example from each cluster at 50% identity, following the instructions provided in their codebase. We did not perform any clustering before training. Because our training set size was similar in size to that used in the original model, we used the recommended 7000 epochs of training. We used the default script with triplet margin loss. After, we performed inference but added our own code into the original CLEAN code to output the classification results into a format compatible with our downstream analysis.
Pika. Pika is a finetuned LLM, which accepts protein sequence as an auxiliary input and can perform reasoning and protein function prediction [66]. We retrained Pika using the training set and evaluated it with the "qa" functionality by creating a new annotation set. The annotation set is an entry for each enzyme with the question: "What is the EC number of this protein?" and the answer of the EC number for that protein. For training, we use 70% of the training data for training, and 15% for testing and validation respectively (the "test" datasets are completely held out of this process). The model was trained for a maximum of 1000 epochs, with a maximum batch size of 100. We used the default models from Pika: esm2_t6_8M_UR50D as the protein model and gpt2 as the text model. All parameters were unchanged from the default apart from increasing the batch size and number of epochs, though hyperparameter tuning could improve the model. To evaluate the ability of Pika to infer the EC numbers, we then query the trained model for each enzyme in the test set with the question: "What is the EC number of this protein?".
CLIPZyme. CLIPZyme is a multimodal contrastive model, between reactions represented as 2D graphs and protein structures represented as 3D graphs [78]. CLIPZyme was trained on direct reaction-protein pairs from the EnzymeMap dataset [35]. In this study, we retrained CLIPZyme model using default parameters and the EGNN protein structure representation for retrieval on the Task 2 train-test splits. We modified the CLIPZyme code to loop over EC numbers and randomly sample pairs of protein sequences and reaction pairs associated with that EC number, much like CREEP, for 40 epochs. In this case, we only provided one protein sequence from each cluster at 50% identity, and we only trained with mapped reactions that had conserved atom numbers on both sides of the reaction, thus reducing the total number of available EC numbers. Furthermore, we dropped several protein sequences with structures that did not work with the EGNN encoder. In the future, the hyperparameters and training procedure could likely be optimized to yield improved performance.
@@ A.5 Extended Benchmarking Results
A.6 License